OpenStudy (anonymous):

how do you solve 4x^(2/3)+16x^(1/3)=-15

OpenStudy (anonymous):

try to move the 4 and the 16 to the other side of the = sign, for a start

OpenStudy (anonymous):

try to move the 4 and the 16 to the other side of the = sign, for a start

OpenStudy (anonymous):

actually thats a good question, i dont know how to do it. anyone?

OpenStudy (anonymous):

I got it, saw this in chat but idk why i cant join the chat. Anyone interested?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

you let a=x^(1/3) so equation becomes 4a^2+16a+15=0 solve that get a=-3/2 or a=-5/2 case 1: a=-3/2 so x^(1/3)=-3/2 x=(-3/2)^3 case 2: similar

OpenStudy (anonymous):

ok, but can you put a step in going from your 2nd line to the third? cause thats also getting me stuck

OpenStudy (anonymous):

ok x^(2/3)=(x^(1/3))^2=a^2 4a^2+16a+15 =4a^2+6a+10a+15 =2a(2a^2+3) + 5(2a+3) =(2a^2+3)(2a+5) hope that's clear enough